# A parallel domain decomposition method for solving elliptic equations on manifolds

### v1

@@ 1 Introduction
Elliptic equations on Riemannian manifolds are important both in analysis and in geometry (see e.g. [46] and [34]). These equations appear in many areas, such as multifluid dynamics, micromagnetics, and image processing (see e.g. [3, 38, 45, 17, 30, 29, 6, 2, 32, 33, 4]). A simple and important example of such equations is
$-\Delta u+bu=f.$ (1)
@@
The idea of [42] was developed by Cao–Qin in [10]. They extended a DDM proposed by P. L. Lions ([36, Section I. 4]) for solving problems on Euclidean domains to a DDM on manifolds. In this paper, we combine the ideas of [42] and [37], which yields an adaption and generalization of Lions’ another DDM ([37, p. 66]). This DDM was also proposed by Lions for solving continuous problems on Euclidean domains. We found that it can also be well-adapted to a DDM on manifolds. One notable advantage of this method is its high level of parallelism, making it particularly valuable when dealing with a large number of subdomains.
Here is an overview of the method in this paper. First of all, global triangulation/grid is completely avoided. There is also no requirement of any compatibility among local grids, i.e. grids over subdomains. (As explained in the introduction of [10], this incompatibility actually reflects the high flexibility of the numerical methods.) The numerical difficulty caused by the nonmatching grids is easily resolved by the technique of interpolation. Furthermore, the subproblems on subdomains can be localized and then be reduced to problems on Euclidean domains. These features bear resemblance to that method described in [10]. However, this paper is remarkably different from [10] in that the current method is highly parallel. In fact, this work follows [37], while [10] follows [36]. Lions described the work [36, Section I. 4] as “sequential” (see the sentence right before (26) in [36]), which means the subproblems on subdomains have to be solved one by one sequentially. On the contrary, the work [37] was claimed as “parallel” (see a few sentences before (37) in [37]), which means the subproblems on subdomains can be solved independently and parallelly at each iterative step.
It is worth noting that, even in the special case where the target manifold is a Euclidean domain, our algorithm (see Algorithms 1 and 2) is different from [37, p. 66] though we follow it largely. Typically, we employ a technique of partition of unity which was not utilized in [37, p. 66]. Due to this technique, we obtain a better convergence theory (see Theorem 2.1 and Remark 2.1).
@@
Our DDM in this paper is numerically verified on examples of $4$-dimensional manifolds. They are $\mathbb{CP}^{2}$, $S^{2}\times S^{2}$, $B^{4}$ and $B^{2}\times S^{2}$, where $\mathbb{CP}^{2}$ is the complex projective plane, $S^{2}$ is the $2$-dimensional unit sphere, and $B^{d}$ is the $d$-dimensional unit ball. The former two manifolds are without boundary, while the latter two are with boundary. The numerical results show that our method solves effectively the equation (1) on these manifolds.
The outline of this paper is as follows. In Section 2, we shall adapt and extend P. L. Lions’ method in [37, p. 66] to a DDM for continuous problems on manifolds and prove its convergence. In Section 3, we shall propose our numerical DDM which is a discrete imitation of Lions’ method. In Section 4, we shall introduce a technique to decompose manifolds with boundary. Finally, some numerical results will be presented in Section 5.
@@ 2 Theory on Continuous Problems
In this section, we shall firstly formulate an elliptic model problem of second order on manifolds. Then we introduce a domain decomposition method (DDM) to solve the model problem (see Algorithm 1 below). This DDM iteration procedure was originally proposed by P. L. Lions in [37, p. 66] for solving differential equations in a Euclidean domain. Unlike those DDMs in [36] and [10], this iterative procedure is highly parallel. More precisely, the subproblem in each subdomain can be solved independently at each iterative step. This procedure motivates us to propose a highly parallel numerical procedure (Algorithm 2) in the next section which provides numerical approximations to the solution of the model problem.
@@ 2.1 Model Problem and Continuous Algorithm
$g=\sum_{\alpha,\beta=1}^{d}g_{\alpha\beta}\mathrm{d}x_{\alpha}\otimes\mathrm{d}x_{\beta},$ (2)
@@
$\Delta u=\frac{1}{\sqrt{G}}\sum_{\alpha=1}^{d}\frac{\partial}{\partial x_{\alpha}}\left(\sum_{\beta=1}^{d}g^{\alpha\beta}\sqrt{G}\frac{\partial u}{\partial x_{\beta}}\right),$
@@
$\left\{\begin{array}[]{rcl}Lu=-\Delta u+bu&=&f,\\ u|_{\partial M}&=&\varphi,\end{array}\right.$ (3)
where $b\geq 0$ is a constant and $f\in L^{2}(M)$. It suffices to solve (3) on each component of $M$. Therefore, without loss of generality, $M$ is assumed to be connected. Furthermore, it is possible for $\partial M$ to be either empty or nonempty. In the case where $\partial M\neq\emptyset$, we assume $\varphi\in C^{0}(\partial M)\cap H^{\frac{1}{2}}(\partial M)$. On the other hand, if $\partial M=\emptyset$, the boundary condition $u|_{\partial M}=\varphi$ is vacuously satisfied. However, to guarantee (3) is well-posed, we need to further assume $b>0$ if $\partial M=\emptyset$. With these assumptions, the problem (3) always has a unique solution $u\in C^{0}(M)\cap H^{1}(M)$.
Now we describe a domain decomposition iteration procedure to solve (3). This method was originally proposed by P. L. Lions in [37, p. 66] for solving differential equations in a Euclidean domain. We found that the method is well-adapted to solve problems on manifolds. Suppose $M$ is decomposed into $m$ subdomains, i.e. $M=\bigcup_{i=1}^{m}M_{i}$. Here $M_{i}$ is a closed subdomain (submanifold with codimension 00) of $M$ with Lipschitz boundary. Let $\gamma_{i}=\partial M_{i}\setminus\partial M$. We make the following assumption on decomposition.
@@
The requirement of the Hausdorff measure of $S$ can be easily satisfied. For example, we may choose $M_{i}$ with piecewise $C^{1}$ boundary and arrange that all $\gamma_{i}$ meet transversely with each other. Transversality is a generic property which, therefore, can be satisfied almost trivially. (For a detailed theory on transversality, see e.g. [28, Chapter 3].)
@@
Unlike those methods in [36, Section I. 4] and [10], Algorithm 1 is highly parallel. Obviously, the subproblems in (4) are solved independently on $M_{i}$ for all $i$. It’s necessary to point out that $u^{n}_{i}$ in (4) is merely defined on $M_{i}$. However, $u^{n}$ is well-defined on the whole $M$ since $\mathrm{supp}\rho_{i}\subset M_{i}$. (More precisely, the function $\rho_{i}u^{n}_{i}$ on $M_{i}$ is extended as 00 outside of $M_{i}$.) We shall prove that $\{u^{n}_{i}\}$ and $\{u^{n}\}$ converge to the solution $u$ in certain sense.
The boundary value condition of $u^{n}_{i}$ in Algorithm 1 is slightly different from Lions’ original approach in [37]. More precisely, a partition of unity was not employed in [37], instead, the boundary value $u^{n}_{i}|_{\gamma_{i}}$ was chosen as an arbitrary function $v$ such that (see [37, (42) & (43)])
@@
Our $u^{n}_{i}|_{\gamma_{i}}$ in Algorithm 1 certainly satisfies the above inequalities. Moreover, by a partition of unity satisfying Assumption 2.1, the iteration in Algorithm 1 even has a better theory of convergence.
@@
Theorem 2.1 was essentially proved by Lions in [37, p. 66-67] in the case that $M$ is a Euclidean domain. Even in that special case, the proof was complicated. Actually, the proof in [37] was presented in a dense style. We shall follow and adapt Lions’ argument and then extend it to the case of general manifolds. Our proof will be given in the next two subsections. To ensure the readability for the reader, we shall provide plenty of details of argument.
@@
Even in the case that $M$ is a Euclidean domain, the statement on convergence in [37] is slightly different from Theorem 2.1. In [37], the function $u^{n}_{i}$ was not necessarily continuous on $M_{i}$, the uniform convergence and $H^{1}$-convergence of $u^{n}_{i}$ on $M_{i}$ were not claimed either. However, it was proved that $u^{n}_{i}$ converges uniformly to $u$ on each compact set $K_{i}\subset M_{i}\setminus\overline{\gamma_{i}}$ there.
Remark 2.7.
One may obtain a convergence theory better than Theorem 2.1 if one imposes an additional assumption as follows: every three distinct subdomains have an empty intersection (cf. [37, (36)]). As pointed in [37, Proposition 2 and p. 63-65], under this assumption, one can prove that the convergence is geometric, which is stronger than the conclusion of Theorem 2.1, by an alternative and much easier approach. However, this additional assumption is stringent in practice. In fact, none of our numerical examples satisfies this assumption.
@@ 2.2 Subsolutions and Supersolutions
$\left\{\begin{array}[]{rcl}L\hat{u}&=&0,\\ \hat{u}|_{\partial M}&=&0,\end{array}\right.$ (5)
@@
$Lu=-\frac{1}{\sqrt{G}}\sum_{\alpha=1}^{d}\frac{\partial}{\partial x_{\alpha}}\left(\sum_{\beta=1}^{d}g^{\alpha\beta}\sqrt{G}\frac{\partial u}{\partial x_{\beta}}\right)+bu$ (6)
@@
There is also a notion of subsolutions and supersolutions in the sense of distribution. It is actually equivalent to that in the sense of viscosity (see e.g. [35, Appendix], [31], and [26]).
@@
$\partial M_{i}\setminus S=\partial M_{i}\setminus(\gamma_{i}\cap S)=(\partial M_{i}\cap\partial M)\cup(\gamma_{i}\setminus S).$ (9)
@@
We firstly claim $\underline{u}_{i}|_{\partial M_{i}}$ is continuous at each $x\in\gamma_{i}\setminus S$. (See Fig. 2 for an illustration.) Suppose $y\in\partial M_{i}$. Since $x\notin S$, we see $y\notin\partial M\cup S$ when $y$ is close enough to $x$. In other words, $y\notin\partial M$ and $y\notin\partial M_{j}$ for all $j\neq i$. Thus, when $y\in\partial M_{i}$ is near $x$, we have $y$ belongs to the same subdomains $M_{j}$ as $x$ does and
@@
[…]
@@
$\{\underline{u}^{n}_{i}\}$ and $\{\bar{u}^{n}_{i}\}$ converge to 00 on $M_{i}\setminus(\gamma_{i}\cap S)$. The convergence is uniform on each $K\subset\subset M_{i}\setminus(\gamma_{i}\cap S)$.
Proof 2.15.
@@ 2.3 Proof of Convergence
Proof 2.20 (Proof of Theorem 2.1).
@@ 3.1 Finite Element Spaces over a d-Rectangle
Then $[a_{i},b_{i}]$ is divided into $N_{i}$ subintervals. The refinement of all such $[a_{i},b_{i}]$ provides a grid on $D$. This divides $D$ into $\prod_{i=1}^{d}N_{i}$ many small $d$-rectangles
$\prod_{i=1}^{d}[c_{i,t_{i}-1},c_{i,t_{i}}],$ (14)
where $1\leq t_{i}\leq N_{i}$ for all $i$. Each small $d$-rectangle (14) is an element of the grid. We refer to [10, § 3.1] (see also [11, p. 56-64]) for a detailed description of the finite element space of $d$-rectangles of type ($1$) based on such a grid.

### v2

@@ 1 Introduction
Elliptic equations on Riemannian manifolds are important both in analysis and in geometry (see e.g. [47] and [34]). These equations appear in many areas, such as multifluid dynamics, micromagnetics, and image processing (see e.g. [3, 39, 46, 17, 30, 29, 6, 2, 32, 33, 4]). Recently, numerical methods for high dimensional spheres are developed using machine learning techniques (see e.g. [35]). A simple and important example of such equations is
(1) $-\Delta u+bu=f.$
@@
The idea of [43] was developed by Cao–Qin in [10]. They extended a DDM proposed by P. L. Lions ([37, Section I. 4]) for solving problems on Euclidean domains to a DDM on manifolds. In this paper, we combine the ideas of [43] and [38], which yields an adaption and generalization of Lions’ another DDM ([38, p. 66]). This DDM was also proposed by Lions for solving continuous problems on Euclidean domains. We found that it can also be well-adapted to a DDM on manifolds. One notable advantage of this method is its high level of parallelism, making it particularly valuable when dealing with a large number of subdomains.
Here is an overview of the method in this paper. First of all, global triangulation/grid is completely avoided. There is also no requirement of any compatibility among local grids, i.e. grids over subdomains. (As explained in the introduction of [10], this incompatibility actually reflects the high flexibility of the numerical methods.) The numerical difficulty caused by the nonmatching grids is resolved by the technique of interpolation. Furthermore, the subproblems on subdomains can be localized and then be reduced to problems on Euclidean domains. These features bear resemblance to that method described in [10]. However, this paper is remarkably different from [10] in that the current method is highly parallel. In fact, this work follows [38], while [10] follows [37]. Lions described the work [37, Section I. 4] as “sequential” (see the sentence right before (26) in [37]), which means the subproblems on subdomains have to be solved one by one sequentially. On the contrary, the work [38] was claimed as “parallel” (see a few sentences before (37) in [38]), which means the subproblems on subdomains can be solved independently and parallelly at each iterative step.
It is worth noting that, even in the special case where the target manifold is a Euclidean domain, our algorithm (see Algorithms 1 and 2) is different from [38, p. 66] though we follow it largely. Typically, we employ a technique of partition of unity which was not utilized in [38, p. 66]. Due to this technique, we obtain a better convergence theory (see Theorem 2.3 and Remark 2.4).
@@
Our DDM in this paper is numerically verified on examples of $4$-dimensional manifolds. They are $\mathbb{CP}^{2}$, $B^{4}$ and $B^{2}\times S^{2}$, where $\mathbb{CP}^{2}$ is the complex projective plane, $S^{2}$ is the $2$-dimensional unit sphere, and $B^{d}$ is the $d$-dimensional unit ball. The $\mathbb{CP}^{2}$ is without boundary, while the latter two are with boundary. The numerical results show that our method solves effectively the equation (1) on these manifolds.
The outline of this paper is as follows. In Section 2, we shall adapt and extend P. L. Lions’ method in [38, p. 66] to a DDM for continuous problems on manifolds and prove its convergence. In Section 3, we shall propose our numerical DDM which is a discrete imitation of Lions’ method. In Section 4, we shall introduce a technique to decompose manifolds with boundary. Finally, some numerical results will be presented in Section 5.
@@ 2 Theory on Continuous Problems
In this section, we shall firstly formulate an elliptic model problem of second order on manifolds. Then we introduce a domain decomposition method (DDM) to solve the model problem (see Algorithm 1 below). This DDM iteration procedure was originally proposed by P. L. Lions in [38, p. 66] for solving differential equations in a Euclidean domain. Unlike those DDMs in [37] and [10], this iterative procedure is highly parallel. More precisely, the subproblem in each subdomain can be solved independently at each iterative step. This procedure motivates us to propose a highly parallel numerical procedure (Algorithm 2) in the next section which provides numerical approximations to the solution to the model problem.
@@ 2.1 Model Problem and Continuous Algorithm
(2) $g=\sum_{\alpha,\beta=1}^{d}g_{\alpha\beta}\mathrm{d}x_{\alpha}\otimes\mathrm{d}x_{\beta},$
@@
$\Delta u$ $=\frac{1}{\sqrt{G}}\sum_{\alpha=1}^{d}\frac{\partial}{\partial x_{\alpha}}\left(\sum_{\beta=1}^{d}g^{\alpha\beta}\sqrt{G}\frac{\partial u}{\partial x_{\beta}}\right)$
(3) $=\sum_{\alpha,\beta=1}^{d}g^{\alpha\beta}\frac{\partial^{2}u}{\partial x_{\alpha}\partial x_{\beta}}+\frac{1}{\sqrt{G}}\sum_{\beta=1}^{d}\sum_{\alpha=1}^{d}\frac{\partial}{\partial x_{\alpha}}\left(g^{\alpha\beta}\sqrt{G}\right)\frac{\partial u}{\partial x_{\beta}},$
@@
(4) $\left\{\begin{array}[]{rcl}Lu=-\Delta u+bu&=&f,\\ u|_{\partial M}&=&\varphi,\end{array}\right.$
where $b\geq 0$ is a constant and $f\in L^{2}(M)$. It suffices to solve (4) on each component of $M$. Therefore, without loss of generality, $M$ is assumed to be connected. Furthermore, it is possible for $\partial M$ to be either empty or nonempty. In the case where $\partial M\neq\emptyset$, we assume $\varphi\in C^{0}(\partial M)\cap H^{\frac{1}{2}}(\partial M)$. On the other hand, if $\partial M=\emptyset$, the boundary condition $u|_{\partial M}=\varphi$ is vacuously satisfied. However, to guarantee (4) is well-posed, we need to further assume $b>0$ if $\partial M=\emptyset$. With these assumptions, the problem (4) always has a unique solution $u\in C^{0}(M)\cap H^{1}(M)$. By (2.1), the local problems of (4) are general elliptic equations with varying coefficients.
Now we describe a domain decomposition iteration procedure to solve (4). This method was originally proposed by P. L. Lions in [38, p. 66] for solving differential equations in a Euclidean domain. We found that the method is well-adapted to solve problems on manifolds. Suppose $M$ is decomposed into $m$ subdomains, i.e. $M=\bigcup_{i=1}^{m}M_{i}$. Here $M_{i}$ is a closed subdomain (submanifold with codimension 00) of $M$ with Lipschitz boundary. Let $\gamma_{i}=\partial M_{i}\setminus\partial M$. We make the following assumption on decomposition.
@@
The requirement of the Hausdorff measure of $S$ can be easily satisfied. For example, we may choose $M_{i}$ with piecewise $C^{1}$ boundary and arrange that all $\gamma_{i}$ meet transversely with each other. Transversality is a generic property which, therefore, can be satisfied almost trivially. (For a detailed theory on transversality, see e.g. [28, Chapter 3].) However, a nonoverlapping decomposition, commonly used in DDMs, does not satisfy Assumption 2.1. Our DDM is an overlapping one.
@@
Unlike those methods in [37, Section I. 4] and [10], Algorithm 1 is highly parallel. Obviously, the subproblems in (5) are solved independently on $M_{i}$ for all $i$. It’s necessary to point out that $u^{n}_{i}$ in (5) is merely defined on $M_{i}$. However, $u^{n}$ is well-defined on the whole $M$ since $\mathrm{supp}\rho_{i}\subset M_{i}$. (More precisely, the function $\rho_{i}u^{n}_{i}$ on $M_{i}$ is extended as 00 outside of $M_{i}$.) We shall prove that $\{u^{n}_{i}\}$ and $\{u^{n}\}$ converge to the solution $u$ in certain sense.
The boundary value condition of $u^{n}_{i}$ in Algorithm 1 is slightly different from Lions’ original approach in [38]. More precisely, a partition of unity was not employed in [38], instead, the boundary value $u^{n}_{i}|_{\gamma_{i}}$ was chosen as an arbitrary function $v$ such that (see [38, (42) & (43)])
@@
Our $u^{n}_{i}|_{\gamma_{i}}$ in Algorithm 1 certainly satisfies the above inequalities. Moreover, by a partition of unity satisfying Assumption 2.1, the iteration in Algorithm 1 even has a better theory of convergence, see Theorem 2.3 and Remark 2.4 below.
@@
Theorem 2.3 was essentially proved by Lions in [38, p. 66-67] in the case that $M$ is a Euclidean domain. Even in that special case, the proof was complicated. Actually, the proof in [38] was presented in a dense style. We shall follow and adapt Lions’ argument and then extend it to the case of general manifolds. Our proof will be given in the next two subsections. To ensure the readability for the reader, we shall provide plenty of details of argument.
@@
Even in the case that $M$ is a Euclidean domain, the statement on convergence in [38] is slightly different from Theorem 2.3. In [38], the function $u^{n}_{i}$ was not necessarily continuous on $M_{i}$, the uniform convergence and $H^{1}$-convergence of $u^{n}_{i}$ on $M_{i}$ were not claimed either. It was merely proved that $u^{n}_{i}$ converges uniformly to $u$ on each compact set $K_{i}\subset M_{i}\setminus\overline{\gamma_{i}}$ there. On the other hand, our function $u^{n}_{i}$ is continuous on $M_{i}$, we also obtain the uniform convergence and $H^{1}$-convergence of $u^{n}_{i}$ on $M_{i}$. This is mainly because we employ the technique of partition of unity. The benefits of this technique can be seen in the proof of Theorem 2.3.
@@ 2.2 Subsolutions and Supersolutions
(6) $\left\{\begin{array}[]{rcl}L\hat{u}&=&0,\\ \hat{u}|_{\partial M}&=&0,\end{array}\right.$
@@
(7) $Lu=-\frac{1}{\sqrt{G}}\sum_{\alpha=1}^{d}\frac{\partial}{\partial x_{\alpha}}\left(\sum_{\beta=1}^{d}g^{\alpha\beta}\sqrt{G}\frac{\partial u}{\partial x_{\beta}}\right)+bu$
@@
There is also a notion of subsolutions and supersolutions in the sense of distribution. It is actually equivalent to that in the sense of viscosity (see e.g. [36, Appendix], [31], and [26]).
@@
(10) $\partial M_{i}\setminus S=\partial M_{i}\setminus(\gamma_{i}\cap S)=(\partial M_{i}\cap\partial M)\cup(\gamma_{i}\setminus S).$
@@
We firstly claim $\underline{u}_{i}|_{\partial M_{i}}$ is continuous at each $x\in\gamma_{i}\setminus S$. (See Fig. 2 for an illustration.) Suppose $y\in\partial M_{i}$. Since $x\notin\partial M\cup S$, so does $y$ when $y$ is close enough to $x$. In other words, $y\notin\partial M$ and $y\notin\partial M_{j}$ for all $j\neq i$. Thus, when $y\in\partial M_{i}$ is near $x$, we have $y$ belongs to the same subdomains $M_{j}$ as $x$ does and
@@
$|\underline{u}^{N}_{i}(x)-\underline{u}^{N}_{i}(y)|<\epsilon\qquad\text{and}\qquad|\tilde{u}_{i}(x)-\tilde{u}_{i}(y)|<\epsilon$
@@
$\{\underline{u}^{n}_{i}\}$ and $\{\bar{u}^{n}_{i}\}$ converge to 00 on $M_{i}\setminus(\gamma_{i}\cap S)$. The convergence is uniform on each $K\subset\subset M_{i}\setminus(\gamma_{i}\cap S)$. Here $S$ is the set defined in Assumption 2.1.
Proof 2.12.
@@ 2.3 Proof of Convergence
Proof 2.17 (Proof of Theorem 2.3).
@@ 3 Numerical Scheme
Overall, our numerical Algorithm 2 is an imitation of Algorithm 1 whose convergence has been justified by the main Theorem 2.3. However, the convergence theory of the numerical scheme is not established in this paper. The current work has two objectives. First, we provide a numerical scheme and numerical tests to assess Theorem 2.3. Second, we test the feasibility and efficacy of our numerical scheme by numerical experiments.
@@ 3.1 Finite Element Spaces over a d-Rectangle
Then $[a_{i},b_{i}]$ is divided into $N_{i}$ subintervals. Define a function $\varphi_{i,j}$ on $[a_{i},b_{i}]$ for $0\leq j\leq N_{i}$ as
(15) $\varphi_{i,j}(x_{i})=\begin{cases}\frac{x_{i}-c_{i,j-1}}{c_{i,j}-c_{i,j-1}},&x_{i}\in[c_{i,j-1},c_{i,j}];\\ \frac{x_{i}-c_{i,j+1}}{c_{i,j}-c_{i,j+1}},&x_{i}\in[c_{i,j},c_{i,j+1}];\\ 0,&\text{otherwise}.\end{cases}$
Here $\varphi_{i,j}(x_{i})$ is undefined for $x_{i}<c_{j}$ (resp. $x_{i}>c_{j}$) when $j=0$ (resp. $j=N_{i}$). Clearly, $\varphi_{i,j}$ is piecewise linear such that $\varphi_{i,j}(c_{i,j})=1$ and $\varphi_{i,j}(c_{i,t})=0$ for $t\neq j$.
The refinement of all such $[a_{i},b_{i}]$ provides a grid on $D$. This divides $D$ into $\prod_{i=1}^{d}N_{i}$ many small $d$-rectangles
(16) $\prod_{i=1}^{d}[c_{i,t_{i}-1},c_{i,t_{i}}],$
where $1\leq t_{i}\leq N_{i}$ for all $i$. Each small $d$-rectangle (16) is an element of the grid. A vertex of (16) is a node of the grid which is of the form
$\xi=(c_{1,j_{1}},c_{2,j_{2}},\cdots,c_{d,j_{d}}),$
where $0\leq j_{i}\leq N_{i}$ for all $i$. Let $W_{h}$ be the finite element space of $d$-rectangles of type ($1$) (see [11, p. 57]). A base function in $W_{h}$ associated with the node $\xi$ is
$\varphi_{\xi}(x_{1},\cdots,x_{d})=\prod_{i=1}^{d}\varphi_{i,j_{i}}(x_{i}),$
where $\varphi_{i,j_{i}}$ is the one in (15).